# Relightable and Dynamic Gaussian Avatar Reconstruction from Monocular Video

### v1

@@ 3.2. Relightable and Dynamic Gaussian Avatar
Given a $d$-length pose sequence at frame $t$, $\Theta_{t}\in\mathbb{R}^{d\times J\times 3}=\{\theta_{t-d},\ldots,\theta_{t}\}$, we encode the input motion to temporal feature $f_{t}$ and spatial feature $f_{s}$ respectively. Subsequently, we encode the position feature $f_{x}$ through MLPs. To obtain $f_{t}$, we first reshape the input motion as $f_{p}\in\mathbb{R}^{J\times(d\times z)}$, representing the movement of each joint, representing global motion dynamics. This reshaped motion is then fed into a temporal attention layer. Next, we apply a cross-attention mechanism, where the key and value are derived from the output of the temporal attention layer, while $f_{x}$ serves as the query. Similarly, to obtain $f_{s}$, we embed the spatial attention output along with $f_{x}$. Specifically, we compute the joint difference between $\theta_{t}$ and $\theta_{t-1}$, which is then fed into a spatial attention layer to capture local motion dynamics. Finally, we concatenate the two features and feed them into several layers of MLPs to compute the dynamic skinning weight $\mathcal{W}_{\Theta_{t}}$. The details of $\mathcal{D}_{\mathcal{W}}$ are shown in Fig 3. Based on $\mathcal{D}_{\mathcal{W}}$, we obtain the pose-driven transformation matrices $\mathcal{A}$ which comprises rotation $\mathcal{A}_{R}(\Theta_{t})$ and translation $\mathcal{A}_{T}(\Theta_{t})$, given $\Theta_{t}$ as similar to Eq. LABEL:deformation: $\mathcal{A}_{T}(\Theta_{t})=\big[\mathcal{A}_{R}(\Theta_{t});\mathcal{A}_{T}(\Theta_{t})\big]$. The position $x$, rotation $r$, and normal $n$ of each 3D Gaussian are dynamically transformed as follows:
(1) $x_{o}$ $=\mathcal{A}_{R}(\Theta_{t})\cdot(x_{c}+\Delta x(\theta_{t}))+\mathcal{A}_{T}(\Theta_{t}),$
$r_{o}$ $=\mathcal{A}_{R}(\Theta_{t})\cdot(r_{c}+\Delta r(\theta_{t})),$
$\hat{n}_{o}$ $=\mathcal{A}_{R}(\Theta_{t})\cdot n_{c},$
@@
where $L_{i}(x,\omega_{i})$ and $L_{o}(x,\omega_{o})$ are the incident and outgoing radiance at a position $x$ along direction $\omega_{i}$ and $\omega_{o}$, respectively. $L_{i}(x,\omega_{i})$ is computed by the visibility $v$ and a global light $L(\omega_{i})$ at each Gaussian: $L_{i}(x,\omega_{i})=v(x,\omega_{i})L(\omega_{i})$. We parameterize the visibility $v(x,\omega_{i})$ by 3-degree SH coefficients to present a mono channel. Specifically, we implement view ($\omega^{c}_{i}=c-x$)-dependent visibility, modeled as a SH function: $v(x,\omega^{c}_{i})=\sum_{j}v_{j}Y_{j}(\omega^{c}_{i})$. It is computed in canonical space via a lightweight GCN using $x_{c}$ and $\hat{n}_{o}$ to handle pose variation. $L(\omega_{i})$ is an environment light as a learnable light probe in latitude-longitude format $\in\mathbb{R}^{32\times 64\times 1}$. We employ the Disney Bidirectional Reflectance Distribution Function (BRDF) (burley2012physically) $f(\cdot)$, influenced by the albedo $c_{a}$, normal $n$, roughness $\gamma$, and metallic. We manually set the metallic value to zero to simplify the modeling of geometry and appearance attributes. Thus, the PBR can be defined as follows:
@@ 4. Experiments
4.1. Multi-illuminated Dynamic Human Dataset

### v2

@@ 3.2. Relightable and Dynamic Gaussian Avatar
Given a $d$-length pose sequence at frame $t$, $\Theta_{t}\in\mathbb{R}^{d\times J\times 3}=\{\theta_{t-d},\ldots,\theta_{t}\}$, we encode the input motion to temporal feature $f_{t}$ and spatial feature $f_{s}$ respectively. Subsequently, we encode the position feature $f_{x}$ through MLPs. To obtain $f_{t}$, we first reshape the input motion as $f_{p}\in\mathbb{R}^{J\times(d\times z)}$, representing the movement of each joint, representing global motion dynamics. This reshaped motion is then fed into a temporal attention layer. Next, we apply a cross-attention mechanism, where the key and value are derived from the output of the temporal attention layer, while $f_{x}$ serves as the query. Similarly, to obtain $f_{s}$, we embed the spatial attention output along with $f_{x}$. Specifically, we compute the joint difference between $\theta_{t}$ and $\theta_{t-1}$, which is then fed into a spatial attention layer to capture local motion dynamics. Finally, we concatenate the two features and feed them into several layers of MLPs to compute the dynamic skinning weight $\mathcal{W}_{\Theta_{t}}$. The details of $\mathcal{D}_{\mathcal{W}}$ are shown in Fig 3. Based on $\mathcal{D}_{\mathcal{W}}$, we obtain the pose-driven transformation matrices $\mathcal{A}$ which comprises rotation $\mathcal{A}_{R}(\Theta_{t})$ and translation $\mathcal{A}_{T}(\Theta_{t})$, given $\Theta_{t}$ as similar to Eq. 10: $\mathcal{A}_{T}(\Theta_{t})=\big[\mathcal{A}_{R}(\Theta_{t});\mathcal{A}_{T}(\Theta_{t})\big]$. The position $x$, rotation $r$, and normal $n$ of each 3D Gaussian are dynamically transformed as follows:
(1) $x^{\prime}$ $=\mathcal{A}_{R}(\Theta_{t})\cdot(x_{c}+\Delta x(\theta_{t}))+\mathcal{A}_{T}(\Theta_{t}),$
$r^{\prime}$ $=\mathcal{A}_{R}(\Theta_{t})\cdot(r_{c}+\Delta r(\theta_{t})),$
$\hat{n}$ $=\mathcal{A}_{R}(\Theta_{t})\cdot n_{c},$
@@
where $L_{i}(x,\omega_{i})$ and $L_{o}(x,\omega_{o})$ are the incident and outgoing radiance at a position $x$ along direction $\omega_{i}$ and $\omega_{o}$, respectively. $L_{i}(x,\omega_{i})$ is computed by the visibility $v$ and a global light $L(\omega_{i})$ at each Gaussian: $L_{i}(x,\omega_{i})=v(x,\omega_{i})L(\omega_{i})$. We parameterize the visibility $v(x,\omega_{i})$ by 3-degree SH coefficients to present a mono channel. Specifically, we implement view ($\omega^{c}_{i}=c-x$)-dependent visibility, modeled as a SH function: $v(x,\omega^{c}_{i})=\sum_{j}v_{j}Y_{j}(\omega^{c}_{i})$. It is computed in canonical space via a lightweight MLP using $x_{c}$ and $\hat{n}_{o}$ to handle pose variation. $L(\omega_{i})$ is an environment light as a learnable light probe in latitude-longitude format $\in\mathbb{R}^{32\times 64\times 1}$. We employ the Disney Bidirectional Reflectance Distribution Function (BRDF) (burley2012physically) $f(\cdot)$, influenced by the albedo $c_{a}$, normal $n$, roughness $\gamma$, and metallic. We manually set the metallic value to zero to simplify the modeling of geometry and appearance attributes. Thus, the PBR can be defined as follows:
@@ 4. Experiments
4.1. Multi-view Multi-illuminated Dataset
@@ Acknowledgements.
In this supplementary, we provide more detailed descriptions and experimental results of our proposed framework.
Appendix A Appendix
Figure 10. Example of visible region between $Cam_{a}$ and $Cam_{b}$.
Figure 11. Examples of our constructed database.
A.1. Preliminary
Animatable Avatar Modeling. The human avatar modeling based on NeRF or 3DGS methods typically employ skeleton-driven deformation using a parametric human mesh (e.g., SMPL or SMPL-X). Given skinning weights $\mathcal{W}$ and joint transformation matrices $\{\theta_{k}\}^{J}_{k=1}$, where $J$ is the number of joints, the deformation is performed using the linear blend skinning (LBS) mechanism. To do this, both approaches train the skinning weights for the deformation process.
Specifically, NeRF-based methods learn the skinning weights to transform points from the observation space to the canonical space (inverse skinning approach), whereas 3DGS-based methods regress skinning weights to transform points from the canonical space to the observation space (forward skinning approach). As a result, a point $x_{c}$ in the cannonical space is transformed into a point $x_{o}$ the observation space as follows:
(10) $x_{o}=\mathcal{A}(\theta_{t})\cdot x_{c}=\big(\sum^{J}_{k=1}\mathcal{W}_{k}\theta_{k}\big)\cdot x_{c},$
where, $\mathcal{A}(\theta)$ is a transformation matrices $\big[\mathcal{A}_{R}(\theta_{t});\mathcal{A}_{T}(\theta_{t})\big]$. Additionally, in 3DGS-based methods, Gaussian attributes, such as the rotation $r_{c}$ in the canonical space, are transformed into the observation space as $r_{o}=\mathcal{A}_{R}(\theta_{t})r_{c}$.We note that the skinning weight in the deformation process is fixed weights.
3D Gaussian Splatting. 3D Gaussian Splatting (3DGS) explicitly represents 3D scenes by leveraging a set of 3D Gaussians, which are rendered through a rasterization process. To this end, a 3D Gaussian can be formulated as:
(11) $G(x)=e^{-\frac{1}{2}(x-\mu)^{T}\Sigma^{-1}(x-\mu)},$
where $\mu$ is mean, and $\Sigma$ represents 3D covariance matrix. To ensure the positive semi-definiteness of $\Sigma$, $\Sigma$ is decomposed into quaternion vector $r\in\mathbb{R}^{4}$ and scale vector $s\in\mathbb{R}^{3}$. $r$ and $s$ are converted into a rotation matrix $R$, and a scaling matrix $S$, respectively. By using two matrices, $\Sigma$ is defined as $\Sigma=RSS^{T}R^{T}$. The 3D Gaussians are projected onto the image plane through the splatting process to render the scene from a specific viewpoint. This requires a 2D covariance matrix in the image plane, which can be approximated using the 3D covariance matrix and the projection matrix.
(12) $\Sigma^{\prime}=JW\Sigma W^{T}J^{T},$
where $W$ is a world-to-camera transformation matrix. $J$ represents an approximated projective transformation of Gaussian points. After projection, pixel colors are obtained through alpha blending. Specifically, we count the 2D Gaussians that overlap with each pixel and blend them as follows:
(13) $C=\sum_{i\in N}c_{i}\alpha_{i}\prod^{i-1}_{j=1}(1-\alpha_{i}),$
where $c_{i}$, $\alpha_{i}$ are the color and density of $i$-th 2D Gaussian, respectively.
Method Novel view Novel pose TR
[…]
3DGS-Avatar(qian20243dgs) 30.28 0.9683 0.0317 - - 30.12 0.9567 0.0352 - - 2h
GauHuman(hu2024gauhuman) 31.34 0.9651 0.0305 - - 30.26 0.9516 0.0379 - - 1h
GoMAvatar(wen2024gomavatar) 30.37 0.9689 0.0325 22.57 0.8035 30.34 0.9688 0.0329 20.13 0.758 12h
iHuman(paudel2024ihuman) 28.21 0.9215 0.0342 19.81 0.7112 24.84 0.9067 0.0402 17.65 0.6248 2h
ExAvatar(moon2024expressive) 31.42 0.9588 0.0171 - - 31.54 0.9414 0.0284 - - 2h
EVA(hu2024expressive) 31.65 0.9514 0.0267 - - 30.15 0.9365 0.0317 - - 6h
Ours 33.85 0.9848 0.0115 29.58 0.9671 31.85 0.9748 0.0145 30.58 0.9606 6h
A.2. Detail of Multi-view Multi-illuminated Dataset
The key feature of our MvMi dataset is to capture human performance in high resolution under a wide range of colored lighting conditions. To do this, we established a system with 30 high-resolution cameras ($4096\times 4096$). All cameras are positioned to capture a full $360^{\circ}$ view of a human subject. All cameras are synchronized using external triggers. For various lighting conditions, 19 custom-made LED modules are set up to simulate environmental lighting via Spherical Gaussian (SG) parameters. Each subject was recorded across eight distinct lighting scenarios, with varied poses in each sequence, resulting in approximately 11.5 million high-resolution multi-view video frames. Fig. 11 shows examples from our proposed dataset.
A.3. Training procedure and Inference Pipeline
We present inference pipeline of our method as shown in Fig. We present the details of the training procedure and inference pipeline of our proposed method. In the first stage, our method learns the articulation of avatars based on body motion by training the dynamic skinning weight, which is generated through the pose-dependent weight encoder. Here, $c_{s}$ is utilized for reconstruction guidance. In the second stage, we refine the appearance of avatars by employing PBR process.
A.4. Implementation details
Implementation details of Geometric Consistency Loss. To compute the geometric consistency loss, we need the intersection area on the avatar, which masks the common visible regions between two camera views. In more detail, given two cameras $cam_{a}$ and $cam_{b}$, we can obtain the mask in the viewport of $cam_{a}$ through dot product between the normal and view direction between the avatar and $cam_{b}$. The i-th Gaussian is invisible if $\hat{n}_{i}\cdot(x_{i}-cam_{b})\leq t$. We set $t=9^{\circ}$. We present an example of the masks in Fig 10.
Implementation details of training. Before setting the Gaussians on the vertices of SMPL (smpl), we subsample the vertices to approximately 30K. We do not perform Gaussian cloning, splitting, or pruning as in 3DGS (3dgs). We optimize the total objectives using the Adam optimizer with a learning rate of $1e^{-3}$. The batch size is set to 1, and training is conducted on a single NVIDIA A6000 Ada GPU for 6 hours. We select one camera view from our dataset and use the first 4/5 of its frames as training data. The remaining frames across all camera views are used to evaluate novel pose rendering. For evaluating novel view rendering, we use the corresponding frames from all camera views except the selected view.
Appendix B Additional Results
We conducted additional comparison on ZJU-MoCap database. Quantitative Results We show PSNR, SSIM, LPIPS, PSNRn and SSIMn on the ZJU-MoCap database in Tab 7. As shown in the table, our method achieves high fidelity of both rendering quality and normal compared to the state-of-the-art methods.
Qualitative Results We present a qualitative result on the ZJU-MoCap dataset to demonstrate the reconstruction as shown in Fig. LABEL:Fig:qa_recon1. Our proposed framework can produce photo-realistsic human avatar with fine-grained geoemtry. This ensure the photo-realistsic rendering results under arbitrary lighting environment as shwon in Fig. 12.
Additional Analysis As shown in Tables 2 and 7, we observe that the quantitative performance is slightly lower when trained on our proposed dataset compared to the ZJU dataset. To investigate this discrepancy, we conducted a detailed analysis of the differences between the two datasets. Fig. 13 (a) illustrates the differences in capture environments: the left side shows the camera setup used in the ZJU dataset, while the right side depicts the camera arrangement in our proposed dataset. Notably, the subject-to-camera distance in the ZJU dataset is shorter than in our setup, suggesting that the range of motion appears more compact in the captured images. Additionally, we explicitly analyzed the differences between the two datasets. To ensure a fair comparison, we resized all images from both datasets to a resolution of $512\times 512$ and extracted the subject’s bounding box in each frame. We then measured the width ($\Delta x$) and height ($\Delta y$) of each bounding box and normalized these values by dividing them by 512. The results are shown in Fig. 13 (b). We observe that, in general, the bounding box size in our dataset is larger compared to that in the ZJU dataset, leading to a wide range of variations in the image domain. This difference may affect the visual motion cue, contributing to the performance gap.
Appendix C Discussion
Our proposed framework can reconstruct a high-fidelity human avatar and also enables rendering the avatar under arbitrary lighting conditions using PBR process. However, PBR process is difficult to present complex reflectance. While ray tracing could address this, its high computational complexity led us to use a rough approximation. To mitigate these limitations, we considers the two most important dimensions: (1) a well-constructed dataset, and (2) a well-designed modeling framework. Still, our database can benefit more from more diverse reflectance scenarios, and our framework can be equipped with better generative models such as diffusion models.
Figure 12. Qualitative results of human avatar relighting on ZJU-MoCap database.
Figure 13. Comparison of (a) the capture systems and (b) the bounding box width and height distributions between the ZJU Mocap (peng2021neural) and our proposed dataset.